From the following balanced equation:

3NaOH + H3​PO4​ Na3​PO4​ + 3H2​O

 How many moles of Na3PO4 can be made from 4.002 moles of H3PO4 and excess NaOH?


1
Expert's answer
2021-04-11T23:50:45-0400

Molar Mass of NaOH = 39.997

Molar Mass of H3PO4 = 97.994

Molar Mass of Na3PO4 = 163.9

1/4 × 4.002 = 1.0005 moles of H3PO4

3/4× 4.002 = 3.0015 moles of NaOH

From the equation

1 mole of H3PO4 = 1 mole of Na3PO4

1.0005 moles of Na3PO4

3 moles of NaOH = 1 mole of Na3PO4

= 3.0015/3 = 1.0005

= 1.0005 × 2 = 2.001 moles of Na3PO4